The following probability model gives the probability structure of a lottery ticket sold in Florida.
X -$10 $1 $3 $10 $20
P(X) 0.5 0.25 0.15 0.09 0.01


Calculate the mean, variance, and standard deviation.
1
Expert's answer
2020-12-24T15:57:21-0500

Mean ="\sum X. P(X)"

=(-10*0.5)+(1*0.25)+(3*0.15)+...+(20*0.01)

=-3.2

Variance ="\sum X^2 P(X) - \mu ^2"

"=((-10^2*0.5)+(1^2*0.25)+...+(20^2*0.01))-(-3.2^2)"

=14.36

Standard deviation ="\sqrt {variance}"

="\sqrt {14.36}"

=3.879